### v2

# LLMGA: Multimodal Large Language Model based Generation Assistant

Abstract
In this paper, we introduce a Multimodal Large Language Model-based Generation Assistant (LLMGA), leveraging the vast reservoir of knowledge and proficiency in reasoning, comprehension, and response inherent in Large Language Models (LLMs) to assist users in image generation and editing. Diverging from existing approaches where Multimodal Large Language Models (MLLMs) generate fixed-size embeddings to control Stable Diffusion (SD), our LLMGA provides a detailed language generation prompt for precise control over SD. This not only augments LLM context understanding but also reduces noise in generation prompts, yields images with more intricate and precise content, and elevates the interpretability of the network. To this end, we curate a comprehensive dataset comprising prompt refinement, similar image generation, inpainting &\&& outpainting, and visual question answering. Moreover, we propose a two-stage training scheme. In the first stage, we train the MLLM to grasp the properties of image generation and editing, enabling it to generate detailed prompts. In the second stage, we optimize SD to align with the MLLM’s generation prompts. Additionally, we propose a reference-based restoration network to alleviate texture, brightness, and contrast disparities between generated and preserved regions during image editing. Extensive results show that LLMGA has promising generative capabilities and can enable wider applications in an interactive manner.
1 Introduction
Artificial Intelligence Generated Content (AIGC) has witnessed remarkable advancements, particularly propelled by the evolution of large language models (LLMs) [9, 11, 53] for text generation and diffusion models (DMs) [22, 40] for image generation. LLMs, in particular, have received considerable acclaim for their exceptional ability to comprehend, reason, make decisions, possess extensive knowledge, and generate text with unparalleled accuracy and fluency.
Recent studies have begun delving deeper into Multimodal Large Language Models (MLLMs) [3, 67] built upon LLMs, aiming to empower LLMs to comprehend inputs extending beyond text. For example, BLIP-2 [29] and LLaVA [31] employ visual encoders to transform images into input embeddings, enabling them to be used as prompts alongside text input within the LLM, thus achieving compatibility of LLMs with the visual modality. Furthermore, recent works focused on extending the capabilities of LLMs to generate multimodal outputs. For example, GILL [28] involves instructing LLMs to predict fixed-size visual embeddings aligned with CLIP [38] space to control the Stable Diffusion [40] (SD) for image generation.
Figure 1: Examples of LLMGA for assisting in image generation and editing. (1) For T2I generation, LLMGA can refine the user’s generation prompt to produce more vivid and vibrant images. (2) For similar image generation, LLMGA can understand the component and layout of the input images and generate a similar image. (3) For inpainting and outpainting, LLMGA can provide detailed generation prompts based on user preferences and input images. (4) Interactive image generation and editing exemplify the comprehensive capabilities of LLMGA. Users can design satisfactory images by engaging in interactions with LLMGA, leveraging its vast knowledge and ideas.
However, existing works [28, 50] merely focus on enabling LLM to output images but do not aim to assist users in generating or editing images to enhance quality. In this paper, we aim to develop a Multimodal Large Language Model-based Generation Assistant (LLMGA) to better assist image generation models, making them more user-friendly and capable of producing high-quality images. In contrast to certain methods [28, 50] that leverage MLLMs to predict fixed-size visual embeddings for implicit SD control, our approach is straightforward. We guide the generation of SD using detailed language prompts from MLLM based on five reasons. (1) The embeddings predicted by the MLLM are often filled with noise. This can be filtered out by mapping them to a fixed language domain, enabling precise control of SD. (2) Detailed language prompts can make the network more transparent and interactive, allowing users to understand MLLM’s thoughts for generating images. (3) MLLM is pre-trained on vast textual datasets. Explicit language prompts rather than implicit embeddings are more advantageous for MLLM to generate prompts and comprehend context. (4) Dynamic-sized language prompt facilitates the addition of generation requests during interactions. (5) Training is more simple and efficient.
However, we face several challenges: (1) MLLM may reject the execution of generation instructions due to its nature as a language assistant. (2) MLLM lacks a comprehensive understanding of image generation and editing, and cannot provide an accurate and detailed generation prompt. (3) Determining which part of texts generated by MLLM to guide SD generation. (4) SD’s CLIP encode only $75$ tokens. Additionally, SD is trained on short captions, whereas our LLMGA typically generates detailed prompts exceeding $150$ tokens. This discrepancy poses a challenge for SD in following the detailed prompt of LLMGA.
[…]
Moreover, we have identified noticeable disparities in texture, contrast, and brightness between the newly generated and preserved sections in SD-based image editing, such as inpainting. To address these issues, we propose a Diffusion-based Reference Restoration Network (DiffRIR), based on the structure of the SOTA restoration network DiffIR [59]. Specifically, aside from images generated by SD, we add masked images as reference inputs into DiffRIR. This enables the DiffRIR to refer to the texture, contrast, and brightness of the preserved regions for restoration. Additionally, we introduce perturbations to contrast and brightness during training, making DiffRIR correct contrast and brightness disparities in the images.
As shown Fig. 1, the remarkable capabilities of the LLMGA come to light, leaving us in awe. LLMGA excels in a myriad of functions, elevating image generation and editing, realizing interactive image editing, and delivering astonishing capabilities for image design due to its expansive world knowledge and robust reasoning abilities.
In summary, our contributions are as follows:
  • We introduced LLMGA, a simple, powerful, and generalizable multimodal generation model. Extensive experiments affirm the efficacy of LLMGA in enhancing image generation and editing through its vast knowledge base.
  • We construct a training dataset, including four parts: prompt refinement, similar image generation, inpainting & outpainting, and visual question answering. This enhances LLMGA’s comprehension of generation and editing tasks while standardizing response formats.
  • We proposed a restoration network DiffRIR, which introduces reference images and training perturbations to contrast and brightness. DiffRIR can alleviate texture, contrast, and brightness discrepancies between newly generated and preserved regions for edited images.
  • Open-source. The following assets are released: the generated data, the codebase for data generation and model training, the model checkpoint, and a demo.
2 Related Work
Diffusion Model. Diffusion Models (DMs) [48, 22, 49, 27, 14, 23, 4, 20, 33, 36, 6] have achieved remarkable results in image generation. DMs adopt a parameterized Markov chain to optimize the lower variational bound on the likelihood function. In this way, it can generate realistic images from Gaussian noise. After that, several DM methods [59, 58, 32, 42, 7, 17, 25, 26, 54, 8, 37, 61, 18, 5, 44, 13] have been tailored to enhance the text-to-image (T2I) generation and editing. Notably, GLIDE [34] pioneered the incorporation of text features into transformer blocks during the denoising process. Subsequently, DALL-E [39], Imagen [43], and Stable Diffusion [40] have made substantial strides in improving T2I generation. Subsequently, some works  [64, 41, 65] introduced conditioning controls to the DMs to facilitate a more convenient and precise manipulation of the generation process. Overall, enhancing the user-friendliness of DMs is a key focus within the community. In this paper, we introduce LLMGA, leveraging the extensive knowledge and powerful reasoning capabilities of LLM to facilitate users in achieving more easily attainable and satisfactory image designs.
Multimodal Large Language Models. Recently, LLMs have undeniably made profound impacts and revolutions within the entire AI community and beyond. For example, exemplary LLMs, such as ChatGPT and GPT4 [35], have showcased remarkable abilities in comprehension, reasoning, responses, and knowledge reservoirs. Subsequently, a range of LLMs [12], including Vicuna [10], LLaMA [53], and Alpaca [51] have been released as open-source models, substantially propelling advancements of the community.
Afterward, the community began focusing on the development of the Multimodal Large Language Model [67, 24, 16, 62, 15, 19, 63, 2]. They aim to enable LLMs to comprehend both images and text and provide textual responses. For instance, Flamingo [3] encodes images and feeds them into the LLM’s attention layer. BLIP-2 [29] employs Q-Former to encode input images into queries. Furthermore, LLaVA [31] utilizes CLIP [38] to encode images into image embeddings and concatenate them with text embeddings.
Recent concurrent works, such as Next-GPT [57], have extended the capabilities to encompass audio and video modalities. Moreover, Visual-ChatGPT [56] and HuggingGPT [47] make LLMs act as agents capable of invoking various pre-trained visual models to achieve MLLM. However, these works focus on making LLM determine the combined invocation of modules (such as detection, recognition, and generation) to fulfill user requirements. However, these methods are not tailored for generation and editing and use concise prompts that lack the capability to enhance results. Thus, we propose LLMGA, which is designed to assist with image generation and editing. It can achieve satisfactory results by reasoning and interaction with the user.
3 Methodology
Figure 2: The overview of LLMGA. (a) In the first training stage, we train the MLLM to produce generation prompts based on provided instructions. Moreover, we construct a training dataset including four categories: prompt refinement, similar image generation, inpainting & outpainting, and visual question answering. (b) In the second training stage, we optimize SD to be adapt to the detailed generation prompts from MLLM. (c) In the inference stage, LLMGA can respond to user queries and assist in image generation and editing.
3.1 Overview of LLMGA
In this paper, we aim to design a MLLM-based Generation Assistant (LLMGA). Our LLMGA produces detailed language-based generation prompts to control SD rather than predicting fixed-sized visual embeddings [28, 50] to govern SD. This has five advantages: (1) Visual embeddings contain noise, and mapping them to the language domain can filter out this noise, enabling precise SD control. (2) Language-based generation prompts facilitate users in understanding the LLMGA’s thoughts, enhancing interaction. (3) Dynamic-sized language-based generation prompt enables the addition of generation requests. (4) MLLM is pre-trained on textual datasets. Language prompts rather than implicit visual embeddings are more advantageous for MLLM to generate accurate prompts and comprehend context. (5) Training is simpler and more efficient.
However, we need to address several issues: (1) As a language assistant, MLLM may decline the execution of generation instructions. (2) MLLM lacks a nuanced understanding of image generation and editing, and cannot produce precise and detailed generation prompts. (3) MLLM needs to decide which part of the output text serves as generation prompts to guide generation. (4) SD’s CLIP encode only $75$ tokens. Moreover, SD is trained on short captions, while detailed prompts generated by LLMGA may exceed $150$ tokens.This disparity makes it hard for SD to understand the instructions from LLMGA. To this end, we construct a training dataset and two-stage training schemes, which train the MLLM (Sec. 3.2) and SD (Sec. 3.3).
The network structure and pipeline of LLMGA are illustrated in Fig. 2. Specifically, as shown in Fig. 2 (a) and (c), the images $\mathbf{I}_{input}$ are encoded into image embeddings by CLIP vision encoder and a projection layer. Subsequently, the image embedding is concatenated with the text embedding and fed into the LLM to obtain text output $\mathbf{T}_{output}$. This process can be formulated as:
$\vspace{-1mm}\mathbf{T}_{output}=\mathcal{F}_{MLLM}(\mathbf{T}_{input},\mathbf{I}_{input}),$ (1)
where $\mathbf{T}_{input}$ indicates the input text instructions. Notably, $\mathcal{F}_{MLLM}$ can process only $\mathbf{T}_{input}$ as input.
[…]
Next, we output $\mathbf{T}_{R}$ directly, while $\mathbf{T}_{P}$ is fed into the subsequent SD for T2I generation (Eq. 2) or editing (Eq. 3). Notably, for image editing such as inpainting and outpainting, an additional mask $\mathbf{I}_{mask}$ and the edited image $\mathbf{I}_{input}$ are essential inputs for the inpainting SD. These inputs specify the region requiring editing.
$\vspace{-1mm}\mathbf{I}_{T2I}=\mathcal{F}_{T2I}(\mathbf{T}_{P},\mathbf{Z}),$ (2)
$\vspace{-1mm}\mathbf{I}_{edit}=\mathcal{F}_{edit}(\mathbf{I}_{input},\mathbf{I}_{mask},\mathbf{T}_{P},\mathbf{Z}),$ (3)
Where $\mathbf{Z}$ denotes the random Gaussian noise. Furthermore, to ensure the encoding of all $\mathbf{T}_{P}$ for SD, we iteratively run the CLIP text encoder until all prompts are encoded.
3.2 MLLM Training
As described in Sec. 3.1, original MLLMs are specifically designed and trained as language assistants, but they lack the proficiency to assist in image generation and editing. Notably, considering the input of images and the performance of open-sourced LLMs, employing few-shot learning to guide the model to achieve the desired results proves challenging and inefficient. Therefore, it is crucial to train the MLLM to be a generation assistant, know the response format, and elevate its comprehension of the properties of image generation and editing. Thus, as shown in Fig. 2(a), we construct a training dataset consisting of four categories:
[…]
    Figure 3: Visual comparison on T2I. LLMGA can refine short prompts by adding details, such as clothing, background, and actions.
  • Visual Question Answering. To preserve the Visual Question Answering (VQA) capability of MLLM, we have integrated the VQA dataset into the training of LLMGA. Specifically, we utilized the LLaVA-Instruct-150k dataset [31], which was generated by GPT-4.
As illustrated in Fig. 2 (a), we freeze the CLIP vision encoder and optimize the projection layer and LLM. The model is trained end-to-end using the auto-regressive cross-entropy loss ($\mathcal{L}_{MLLM}$) for text generation. Given the ground-truth targets $\mathbf{T}_{GT}$, this loss can be formulated as:
$\mathcal{L}_{MLLM}=\mathbf{CE}\left(\mathbf{T}_{output},\mathbf{T}_{GT}\right).$ (4)
3.3 Stable Diffusion Training
As described in Sec. 3.1, the original SD’s CLIP text encoder only encodes $75$ tokens, which cannot handle the entire MLLM’s generation prompt. Moreover, the original SD is trained on brief captions, which cannot fully understand the generation prompts. Thus, we repeatedly use the VAE text encoder to encode all instances of $\mathbf{T}_{P}$ for SD. Besides, we train the T2I SD model and the inpainting SD model, respectively. For both generation and editing tasks, the generation prompts of MLLM are detailed descriptions of images. Therefore, during training, we can instruct MLLM to provide a detailed description $\mathbf{T}_{P}$ for images from the LAION-Aesthetics and MSCOCO datasets. Subsequently, $\mathbf{T}_{P}$ is fed into T2I SD or inpainting SD for joint training. Notably, we only optimize the SD unet while freezing the parameters of other networks. To accelerate the training process, we record the prompt $\mathbf{T}_{P}$ of MLLM to avoid redundant calculations. The model is trained using SD loss:
$\vspace{-2mm}\mathcal{L}_{SD}=\mathbb{E}_{\mathbf{Z}_{t},\mathbf{C},\epsilon,t}\left(\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{Z}_{t},\mathbf{C}\right)\right\|_{2}^{2}\right),$ (5)
where $\mathbf{Z}_{t}=\sqrt{\overline{\alpha_{t}}}\mathbf{Z}_{0}+\sqrt{1-\overline{\alpha_{t}}}\epsilon$ represents the noised feature map at timestep $t$. Ground truth images are encoded into latent space to derive $\mathbf{Z}_{0}$. Here, $\epsilon\in\mathcal{N}(0,\mathbf{I})$ represents Gaussian noise, and $\epsilon_{\theta}$ refers to the SD unet. $\mathbf{C}$ indicates the conditional information. For T2I generation, $\mathbf{C}$ is $\mathbf{T}_{P}$. For inpainting and outpainting, $\mathbf{C}$ contains $\mathbf{T}_{P}$, the mask, and the VAE-encoded masked image.
Table 1: Quantitative comparison on T2I generation on the subset of MSCOCO dataset. The best and second best performances are marked in bold and underlined, respectively.
Method FIDnormal-$\downarrow$ ISnormal-$\uparrow$ CLIP-Inormal-$\uparrow$
SD1.5 142.86 13.50 68.93
SD1.5-ft 141.99 13.57 71.05
GILL [28] 150.07 12.19 67.63
LLMGA-7B (SD1.5) 147.13 13.39 70.21
LLMGA-7B (SD1.5-ft, Ours) 139.46 13.63 72.26
LLMGA-13b (SD1.5) 145.97 13.41 70.24
LLMGA-13b (SD1.5-ft, Ours) 136.23 13.69 73.24
Figure 4: Visual comparison on inpainting and outpainting. LLMGA can infer complete images based on input masked images.
3.4 Restoration Network Training
For SD inpainting and outpainting, we observed noticeable disparities between the preserved and newly generated regions in the edited images. These disparities can be attributed to variations in texture, contrast, and brightness. Besides, after finishing image generation and editing, it is customary to employ image restoration methods [59, 58, 55, 60] on the generated images to enhance the quality. Therefore, we intend to employ a restoration network to address these disparities in the edited images.
To enhance the consistency between the newly generated and the preserved regions, we introduce a reference-based restoration scheme. Specifically, existing restoration methods [59, 55, 60] take low-quality (LQ) images as input and produce high-quality (HQ) images, but they often do not leverage the preserved information from the given masked image. Different from these restoration methods, we concatenate the LQ image ${I}_{LQ}$ and the masked image, i.e., $(1-\mathbf{I}_{mask})\mathbf{I}_{GT}$, as inputs and feed it into the restoration model $\mathcal{F}_{R}$. This process can be formulated as:
$\vspace{-2mm}\mathbf{I}_{HQ}=\mathcal{F}_{R}(concat(\mathbf{I}_{LQ},(1-\mathbf{I}_{mask})\mathbf{I}_{GT})).$ (6)
To further mitigate the brightness and contrast disparities, we introduced additional color degradation (i.e., random brightness and contrast disturbance) into the training process of the restoration model, which is formulated as:
$\vspace{-2mm}\mathcal{D}_{2}(\mathbf{x})=c_{1}\mathbf{x}+c_{2},$ (7)
$\vspace{-1mm}\mathbf{I}_{LQ}=\mathcal{D}_{2}(\mathcal{D}_{1}(\mathbf{I}_{GT})),$ (8)
where $\mathcal{D}_{2}$ indicates contrast and brightness disturbance, $\mathbf{x}$ denotes the input image. $c_{1}$ is the contrast gain, randomly varied within the range of $[0.94,1.06]$, while $c_{2}$ is the brightness bias, randomly varied within the range of $[-0.05,0.05]$. $\mathcal{D}_{1}$ represents the real-world degradation process used in restoration model [55, 59, 60]. $\mathbf{I}_{GT}$ is ground truth image, and $\mathbf{I}_{LQ}$ is LQ image. In this paper, we adopt the network structure of the SOTA restoration model DiffIR [59] and apply our schemes to it to obtain our Diffusion-based Reference Restoration Network (DiffRIR).
4 Experiments
4.1 Implementation Details
For the first stage of training, we employ the pretrained LLaVA-7B or LLaVA-13B as the initial MLLM. We utilize AdamW optimizer, setting the learning rate and weight decay to $2\times 10^{-5}$ and 00, respectively. Moreover, we adopt CosineLR as the learning rate scheduler. The batch size per device and epochs are set to $16$ and $6$, respectively. Besides, the training ratios for prompt refinement, similar image generation, inpainting & outpainting, and visual question answering are specified as $0.5$, $1$, $1$, and $1$, respectively.
For the second stage of training, we adopt the Stable Diffusion 1.5 (SD1.5) as the initial image generation or editing model. We train these models with AdamW optimizer, setting the learning rate to $1\times 10^{-5}$. The batch size is set to $32$. We train SD1.5 by $5\times 10^{4}$ iterations.
For the restoration network, we train our DiffRIR on DIV2K [1] and Flickr2K [52] with the same GAN-based loss function as DiffIR. The batch sizes are set to $64$, and the LQ patch sizes are set to $64\times 64$. We use Adam optimizer, setting the learning rate to $2\times 10^{-4}$. We train this model by $4\times 10^{5}$ iterations.
Table 2: Quantitative comparison for outpainting and inpainting on subset of Places datasets [66].
Method Outpainting Inpainting
Narrow Masks Wide Masks Narrow Masks Wide Masks
FIDnormal-$\downarrow$ LPIPSnormal-$\downarrow$ CLIP-Inormal-$\uparrow$ FIDnormal-$\downarrow$ LPIPSnormal-$\downarrow$ CLIP-Inormal-$\uparrow$ FIDnormal-$\downarrow$ LPIPSnormal-$\downarrow$ CLIP-Inormal-$\uparrow$ FIDnormal-$\downarrow$ LPIPSnormal-$\downarrow$ CLIP-Inormal-$\uparrow$
SD1.5 55.34 0.3146 87.39 79.59 0.4460 42.96 13.78 0.0471 98.80 33.40 0.1291 96.03
SD1.5-ft 54.31 0.3127 88.60 79.54 0.4413 43.07 13.68 0.0466 98.81 33.30 0.1289 96.05
LLMGA-7B (SD1.5) 53.36 0.3090 89.07 75.62 0.4374 43.80 13.26 0.0462 98.82 33.87 0.1269 96.11
LLMGA-7B (SD1.5-ft, Ours) 50.96 0.3059 89.72 70.93 0.4307 44.23 12.63 0.0453 98.87 32.53 0.1245 96.30
LLMGA-13B (SD1.5) 52.92 0.3082 89.69 73.07 0.4370 43.98 12.88 0.0461 98.84 33.58 0.1249 96.22
LLMGA-13B (SD1.5, Ours) 50.73 0.3026 90.20 69.74 0.4296 44.60 12.58 0.0450 98.91 32.39 0.1243 96.38
Figure 5: Visual comparison of image restoration methods. Our DiffRIR can alleviate the texture, contrast, and brightness disparities.
Table 3: Quantitative comparison on image restoration for outpainting images on the subset of Places dataset.
Method Reference Color Degradation FIDnormal-$\downarrow$ LPIPSnormal-$\downarrow$
LLMGA-7B 50.96 0.3059
LLMGA+DiffIR [59] 46.10 0.2247
[…]
4.2 Experimental Results
Evaluation on T2I Generation. The results are shown in Tab. 1. Notably, SD1.5 is the original Stable Diffusion 1.5 while SD1.5-ft indicates the finetuned SD1.5 in LLMGA. We also compare our LLMGA with the recently proposed multimodal generative model GILL [28]. (1) In the 1st and 2nd rows of Tab. 1, we input the original short MSCOCO caption into SD15 and SD1.5-ft, respectively. It is evident that SD1.5-ft, trained with detailed generation prompts in LLMGA, achieves a significant improvement in both image quality and content similarity compared to SD1.5. (2) In the 4th and 5th rows of Tab. 1, we use LLMGA-7B to refine the short MSCOCO caption to the detailed generation prompt and send it to the SD1.5 and SD1.5-ft, respectively. Our LLMGA-7B with SD1.5-ft achieves significant 7.67 FID and 2.05 CLIP-I improvements, respectively. This demonstrates the effectiveness of our stage 2 training. (3) When comparing GILL, our LLMGA-7B achieves notable 10.61 FID and 4.63 CLIP-I improvements, underscoring the effectiveness of LLMGA. (4) Comparing the 5th and 7th rows of Tab. 1, LLMGA-13B exhibits better performance than LLMGA-7B due to its superior reasoning ability. In fact, LLMGA-13B can provide more accurate and detailed answers than LLMGA-7B.
The qualitative results are shown in Fig. 3. LLMGA excels in refining short prompts by incorporating details, including clothing, background, and actions to generate visually rich and pleasing images. In contrast, SD1.5 tends to straightforwardly assemble elements from the short prompts. For example, in the first row of Fig. 3, LLMGA showcases the act of singing by incorporating a microphone. In the second row, LLMGA crafts a battle attire for the husky and places a damaged car in the background, depicting engaging scenarios of battling monsters. Furthermore, as depicted in Fig. 1, LLMGA can refine short captions based on user requirements.
Table 4: Datasets comparison. We conducted comparisons of CLIP-I on T2I, outpainting, and inpainting. The ✓signifies the utilization of the complete dataset during training. Conversely, the absence of the ✓indicates a reduction to only $10\%$ of the original datasets.
Method Prompt Refinement Inpainting &Outpainting Similar Image Generation Visual Question Answering T2I Outpainting Inpainting
[…]
Evaluation on Inpainting and Outpainting. The results are shown in Tab. 2. (1) In the 3rd and 4th rows of Tab. 2, we make LLMGA imagine the complete generation prompts for the given masked images, which are then input into the later SD. Notably, our LLMGA-7B (with SD1.5-ft) demonstrates a significant FID and CLIP-I improvement over LLMGA-7B (with SD1.5) in both outpainting and inpainting. This demonstrates the effectiveness of stage 2 training for image editing. (2) When comparing the 1st and 4th rows of Tab. 2, our LLMGA-7B achieves significant improvements of 8.66 FID and 1.27 CLIP-I over the SD1.5 in outpainting under wide masks. (3) Comparing the 4th and 6th rows of Tab. 2, our LLMGA-13B outperforms LLMGA-7B due to its enhanced knowledge reservoir and superior reasoning capabilities.
The qualitative results are shown in Fig.  4. We can see that LLMGA can deduce and imagine complete images based on masked input images. For example, in the 1st row of Fig. 4, LLMGA can infer the presence of a deer model hanging on the wall based on the given environment. In the 3rd and 4th rows of Fig.4, LLMGA’s rational inference about the scene ensures that the SD editing output aligns with the environmental context, rather than producing chaotic results. Furthermore, as illustrated in Fig. 1, LLMGA can imagine complete images based on user requirements and masked images.
Refer to caption Refer to caption
(a) FID Comparison (b) CLIP-I Comparison
Figure 6: Performance comparison of SD control schemes based on detailed language prompt and embedding.
Evaluation on Image Restoration. The results are shown in Tab. 3. For comparisons, we utilize DiffIR [59] and our DiffRIR on the edited image generated by LLMGA. (1) Comparing the 2nd and 3rd rows of Tab. 3, it is evident that introducing a reference scheme can significantly improve restoration performance. (2) When comparing the 3rd and 4th rows of Tab. 3, it can be observed that introducing color degradation helps alleviate the bright and contrast distortion caused by SD. (3) Comparing the 1st and 4th rows of Tab. 3, our DiffRIR yields significant improvement, further validating the effectiveness of DiffRIR.
[…]
Control SD using detailed language prompt or embedding? The results are shown in Fig. 6. We compare two approaches: GILL, which makes LLM estimate a fix-sized embedding to control SD generation, and LLMGA Embedding, a variant of LLMGA where the language prompt is replaced with embedding, undergoing the same training process as LLMGA. The evaluation is conducted on MSCOCO by instructing these methods to generate images multiple times in conversation form. (1) The quality of generated images (Fig. 6 (a)) and the relevance of content (Fig. 6 (b)) in embedding-based methods (i.e., GILL and LLMGA Embedding) deteriorate rapidly as the number of conversation turns increases. In contrast, our LLMGA remains unaffected. This discrepancy arises from the inherent noise present in the embeddings predicted by LLM. As the number of conversation turns rises, these generated embeddings integrate with the preceding conversations, introducing even more noise. This poses challenges for the precise control of SD-generated content. Our LLMGA addresses this issue by mapping the embedding to the fixed language domain, effectively eliminating such noise. (2) Additionally, LLMGA Embedding also outperforms GILL, indicating that the prompt size used to guide SD generation should be adaptive in content, rather than a fixed size.
[…]
5 Conclusion
LLM possesses an extensive reservoir of knowledge and powerful comprehension and reasoning capabilities. In this paper, we introduce a MLLM-based generation assistant (LLMGA), aiming to exploit LLM’s capabilities in an interactive manner to facilitate more efficient and convenient image generation and editing. Compared to relying on LLM to predict a fixed-size embedding to control SD, we employ detailed generation prompts. These prompts prove to be more favorable for enhancing LLM’s contextual comprehension and generating more accurate and rich content. To this end, we develop a two-stage training scheme and curate a dataset, including four parts: prompt refinement, similar image generation, inpainting & outpainting, and visual question answering. For the first stage, we train MLLM to understand the properties of image generation and editing, enabling it to give fitting responses. For the second stage, we optimize the SD unet to adapt to the generation prompt. Moreover, we propose a DM-based reference restoration network (DiffRIR) to mitigate disparities in texture, contrast, and brightness for image editing. Consequently, LLMGA can offer design suggestions and enhance results based on user’s requests during interactions.
[…]
Supplementary Material
The overview of the supplementary materials:
(1) We offer a comprehensive introduction to the creation of the training dataset (Sec. 6).
(2) Additional details of the training and evaluation for LLMGA are elaborated in this part. Furthermore, a detailed training process for Stable Diffusion XL (SDXL) is also included (Sec. 7).
(3) We present more comparison details on LLMGA and LLMGA Embedding, along with corresponding analyses (Sec. 8).
(4) We provide additional examples showcasing the interactive generation and editing capabilities of LLMGA (with SDXL) (Sec. 9).
(5) More visual results are showcased to highlight LLMGA’s performance in T2I generation (Sec. 10).
(6) More visual results are presented for LLMGA on inpainting and outpainting (Sec. 11).
(7) More Visual results effectively demonstrate the prowess of DiffRIR in addressing brightness and contrast disparities between newly generated and retained regions during image editing, along with its ability to enhance texture details (Sec. 12).
6 Data
For the first stage of training, we constructed a training dataset that requires detailed descriptions of images to assist the LLM in better understanding the compositional details of images and supporting image generation and editing. Specifically, we utilized the MSCOCO [30] and LAION-Aesthetics [46] datasets, encompassing rich real-world scenarios and aesthetically pleasing images, respectively. Then, we employed LLaVA to generate detailed and visually compelling descriptions for these datasets. The prompt format for LLaVA to generate a detailed description is as follows: ”The caption of this image is ’ORIGINAL CAPTION’. INSTRUCTIONS.”, where ’ORIGINAL CAPTION’ is a concise caption provided by the dataset, and ’INSTRUCTIONS’ represents the generation instructions randomly selected from Tab. 5. An example is shown in Fig. 7.
Figure 7: An example of generating detailed descriptions for training data.
Notably, the design of such prompts carries significance. As shown in Fig. 8, we conducted a comparison between two distinct prompts. Firstly, let’s scrutinize the original captions provided by the dataset. They are characterized by brevity, insufficient to encapsulate the entire scene but do include proper nouns. Subsequently, we directed LLaVA to generate a detailed description of the given image using the prompt ”Write a detailed description of the given image.” We observed that LLaVA’s image-based description is comprehensive, evoking a vivid sense of the scene, but it may lack certain proper nouns. However, users frequently include proper nouns, which inherently carry a wealth of information. Therefore, to facilitate the MLLM in learning the correlation between these proper nouns and the scene, it is necessary to introduce proper nouns into descriptions. A straightforward and effective approach (the last row of Fig. 8) is to incorporate the original caption of the image into LLaVA’s prompt. This strategy effectively integrates proper nouns into the detailed description.
”What do you see happening in this image?” ”What do you think is going on in this snapshot?” ”Can you elaborate on the elements of the picture provided?” ”Describe the following image.” ”What’s happening in the scene?” ”Analyze the image in a comprehensive and detailed manner.” ”Write a detailed description of the given image.” ”What is this photo about?” ”Explain the visual content of the image in great detail.” ”What are the key elements in this picture?” ”Can you describe the main features of this image for me?”
Table 5: The list of instructions for image description.
Figure 8: Comparison of three caption types: (1) Original captions are included in the existing datasets, which are concise while encompassing abundant proper nouns. (2) LLaVA descriptions based on images are generated by utilizing LLaVA to deliver detailed depictions of images. While offering a more comprehensive overview with elaborate descriptions, they may not effectively convey certain proper nouns. (3) LLaVA descriptions based on images and captions refer to descriptions generated by guiding LLaVA to deliver comprehensive details, considering both images and captions simultaneously. These descriptions encompass both detailed descriptions and proper nouns (marked in red).
7 More Training and Evaluation Details
In addition to the LLMGA (SD1.5) described in the paper, we have also conducted training for LLMGA (SDXL). LLMGA (SDXL) follows a training process and configuration similar to that of LLMGA (SD1.5) but demonstrates even higher-quality generation. Specifically, for the first training stage, we employ the same MLLM as utilized in LLMGA (SD1.5). Regarding MLLM, we utilize the pretrained LLaVA-7B or LLaVA-13B as the initial MLLM. Our optimization approach involves the use of the AdamW optimizer, with the learning rate set at $2\times 10^{-5}$ and weight decay at 00, respectively. Additionally, we adopt the CosineLR learning rate scheduler. The total batch size and number of epochs are configured at $128$ and $6$, respectively. MLLM training is carried out on our constructed datasets as described in the paper. Furthermore, the training ratios for prompt refinement, similar image generation, inpainting & outpainting, and visual question answering are defined as $0.5$, $1$, $1$, and $1$, respectively.
For the second stage of training, we adopt the Stable Diffusion XL (SDXL) as the initial image generation or editing model. We train these models with AdamW optimizer, setting the learning rate to $1\times 10^{-5}$. The total batch size is set to $32$. We train SD1XL by $5\times 10^{4}$ iterations. For both T2I generation and editing, we conduct training on the LAION-Aesthetic dataset, using the generation prompts generated by the first-stage pretrained MLLM as guidance. The input patch sizes are set to $1024\times 1024$. Moreover, similar to SD1.5, we randomly generate masks for SDXL inpainting and outpainting, including box masks, irregular masks, and boundary masks.
To evaluate our LLMGA quantitatively, we employ a diverse set of metrics. For evaluating the quality and diversity of generated images, we utilize the Inception Score (IS)  [45], and Fréchet Inception Distance (FID) [21]. We leverage CLIP-based metrics [38] CLIP-I to assess the content similarity between generated and ground-truth images.
8 More Details on Control Scheme
The form in which to establish a control link between MLLM and SD is a question that requires careful consideration. In this paper, as illustrated in Fig. 9, we explore two control schemes: namely, our adopted language-based generation prompt control (Fig.9 (a)) and visual embedding-based control scheme (Fig.9 (b)). Here, we will provide a detailed overview of the training approach for LLMGA Embedding, followed by a comparison and analysis of these schemes.
For the training of LLMGA Embedding, we employ a comprehensive three-stage training scheme. (1) In the first stage, for $T_{R}$, we use the same auto-regressive cross-entropy loss ($\mathcal{L}_{MLLM}$, Eq. 9) as LLMGA, and for visual embedding, we utilize the visual embedding loss ($\mathcal{L}_{embed}$, Eq. 10). We simultaneously apply the same training settings as LLMGA. (2) In the second stage, we initiate joint optimization of MLLM and SD. Unlike LLMGA, in this phase, we optimize only the projection layer for LLMGA embedding, using the SD loss ($\mathcal{L}_{SD}$, Eq. 11), and then freeze the parameters of Unet and MLLM. (3) In the third training stage, we freeze the parameters of MLLM and projection, and then optimize the SD Unet using the SD loss ($\mathcal{L}_{SD}$, Eq. 11).
$\mathcal{L}_{MLLM}=\mathbf{CE}\left(\mathbf{T}_{R},\mathbf{T}_{GT}\right),$ (9)
$\mathcal{L}_{embed}=\left\|\phi_{proj}(\mathbf{V}_{E})-\phi_{CLIP}(\mathbf{T}_{caption})\right\|_{2}^{2},$ (10)
$\vspace{-1mm}\mathcal{L}_{SD}=\mathbb{E}_{\mathbf{Z}_{t},\mathbf{C},\epsilon,t}\left(\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{Z}_{t},\mathbf{C}\right)\right\|_{2}^{2}\right),$ (11)
where $\mathbf{T}_{R}$ denotes the generated text response, and $\mathbf{T}_{GT}$ represents the ground-truth target. $\mathbf{CE}(.)$ signifies auto-regressive cross-entropy. $\phi_{proj}(.)$ denotes the projection involving three linear layers. $\phi_{CLIP}(.)$ is the CLIP text encoder. $\mathbf{V}_{E}$ stands for visual embedding, and $\mathbf{T}{caption}$ corresponds to the original caption from the datasets. $\mathcal{L}_{SD}$ represents the diffusion loss as described in the paper.
Figure 9: The illustration of LLMGA and LLMGA Embedding. LLMGA uses language generation prompts as a control scheme, while LLMGA Embedding uses visual embedding as a control scheme.
Figure 10: Examples of LLMGA on interactive generation and editing.
Figure 11: Examples of LLMGA on interactive generation and editing.
Figure 12: Examples of LLMGA on interactive generation and editing.
Figure 13: Visual comparison on T2I. LLMGA can refine short prompts by adding details, such as clothing, background, and actions. In addition, for unfamiliar concepts such as a space elevator, LLMGA can utilize its extensive knowledge base from LLM or accurate generation.
Figure 14: Visual comparison on inpainting and outpainting. LLMGA can infer complete images based on input masked images.
Figure 15: Visual comparison of image restoration methods. Our DiffRIR can alleviate the texture, contrast, and brightness disparities.
The results are presented in Fig. 6 in the paper. Despite comprehensive training, we find that the embedding-based approaches still lags behind LLMGA. Moreover, with an increase in the number of dialogue turns, there is a noticeable decrease for embedding-based approaches in both the accuracy and quality of generation. This observation can be well comprehended through the processing mechanism of LLM. Specifically, LLM predicts an embedding based on previous input images and texts (Eq. 12). After that, embedding is refined through a linear layer, categorizing it into a fixed language domain (Eq. 13). This is because predicted embeddings exist in a continuous space, inherently imprecise and filled with noise. For instance, the same embedding, depending on the sampling probabilities, may generate different semantic words, indicating that embeddings are rife with various forms of noise. Mapping embeddings to a fixed language domain effectively eliminates this noise, enabling precise control over SD generation. The decline in performance of embedding-based methods with an increase in conversation turns is attributed to the introduction of additional noise into the predicted embedding as more prior conversation information is incorporated, thereby affecting accuracy.
$\mathbf{E}=\Psi_{body}(\mathbf{I}_{input},\mathbf{T}_{input}),$ (12)
$\mathbf{T}=\Psi_{linear}(\mathbf{E}),$ (13)
where $\mathbf{I}_{input}$ and $\mathbf{T}_{input}$ represent all preceding input images and texts. $\Psi_{body}$ signifies the network body of LLM, producing an embedding $\mathbf{E}$. $\Psi_{linear}$ constitutes the final linear layer in LLM, tasked with categorizing the embedding into a predetermined text domain, resulting in the generation of text $\mathbf{T}$.
In summary, compared with embedding based methods, our LLMGA using detailed language prompts for control generation has the following advantages:
  • The embeddings predicted by the MLLM are often filled with noise. This can be filtered out by mapping them to a fixed language domain, enabling precise control of SD.
  • Detailed language prompts can make the network more transparent and interactive, allowing users to understand MLLM’s thoughts for generating images.
  • MLLM is pre-trained on vast textual datasets. Explicit language prompts rather than implicit embeddings are more advantageous for MLLM to generate prompts and comprehend context.
  • Dynamic-sized language prompt facilitates the addition of generation requests during interactions.
  • Training is more simple and more efficient.
9 More Results on Interactive Generation and Editing
The results are shown in Figs. 1011, and 12. The results presented here were obtained using LLMGA-7b (SDXL-ft). It can be observed that LLMGA, leveraging the understanding, reasoning abilities, and extensive knowledge repository of LLM, effectively assists users in image design in an interactive manner.
10 More Results on T2I Generation
The more T2I generation visual results are shown in Fig. 13. LLMGA can leverage its vast comprehension, reasoning abilities, and knowledge reservoir to generate visuals with more details. Additionally, for less common concepts like a space elevator, LLMGA can accurately generate images based on its extensive knowledge after providing design descriptions. Furthermore, LLMGA refine prompts based on user requirements.
11 More Results on Inpainting and Outpainting
The more inpainting and outpainting visual results are depicted in Fig. 14. LLMGA can harness its extensive comprehension, reasoning abilities, and knowledge reservoir to infer plausible complete images based on given masked images. For instance, in the first row of Fig. 14, LLMGA can deduce the presence of a boat based on the existence of ripples on the water surface. Additionally, as demonstrated in paper Figs. 1 and 12, LLMGA can edit images in accordance with user specifications and masked images.
12 More Results on Image Restoration
The more image restoration results are shown in Fig. 15. Our DiffRIR can alleviate the texture, brightness, and contrast discrepancies, and generate more realistic details.